OpenStudy (anonymous):

Given sin x =-1/2 with x in quadrant 3 and tan y =4/3 with y in quadrant 1, find the exact value of sin(x-y).

OpenStudy (anonymous):

incomplete info

OpenStudy (anonymous):

thats all the question gives me

OpenStudy (anonymous):

I'm stumped. I don't know if x and y are sides of triangle or angles

OpenStudy (anonymous):

\[actually x=\alpha and y=\beta\]

OpenStudy (anonymous):

can't figure out what you just wrote

OpenStudy (anonymous):

\[x=\alpha \] and \[y=\beta\]

OpenStudy (anonymous):

OK, choice of variables made the question confusing. I got it now. Got to go to class will attempt later.

OpenStudy (anonymous):

okay appreciate it!

OpenStudy (anonymous):

Recall sin (A - B) = Sin A cos B - cos A sin B Use sin A= -1/2 in 3rd quadrant to find cos A Use tan B =4/3 in 1st to find sin of B, cos B Use exact values (no calculator, keep square roots etc) to find above